If a number cube is rolled two times, what is the probability of rolling a 1 both times?
Answer by jim_thompson5910(35256)

P(rolling 1 twice) = P(rolling a 1 AND rolling a 1) = P(rolling a 1) * P(rolling a 1) = (1/6)*(1/6) = (1*1)/(6*6) = 1/36